OpenStudy (johnnewman):

What are the solutions to the equation x2 = 64? A. 32 and –32 B. 8 and –8 C. 32 only D. 8 only

pooja195 (pooja195):

Well we would start with square rooting both sides of the equation since we want to isolate and find the x values \[\huge~\rm~\bf~\sqrt{x^2}=\sqrt{64}\]

OpenStudy (johnnewman):

so i know i need to find what will go into 64

pooja195 (pooja195):

Correct ? x ? = 64

OpenStudy (johnnewman):

I think 32 only

pooja195 (pooja195):

Why? Also, does 32 x 32=64?

OpenStudy (johnnewman):

no i got 1024

pooja195 (pooja195):

So we know that 32 isn't the right answer nor is -32 because -32 x -32=1024 so We know the answer can't be A and C So... ? x ? =64

OpenStudy (johnnewman):

d 8 becaues 8x8=64

pooja195 (pooja195):

Ok 8 is correct but also -8 is correct to because -8 x -8=64 a negative times a negative = a positive

OpenStudy (johnnewman):

so b

pooja195 (pooja195):

Correct ^_^

OpenStudy (johnnewman):

thanks pooja u r the best helper

pooja195 (pooja195):

^_^ yw